Solve 6/x+3 + 2/x = 5x-3/x^2-9 with steps please.



Answer by stanbon(75887)

Solve 6/(x+3) + 2/x = 5x-3/(x^2-9) with steps please.
---------------
Multiply thru by x(x^2-9) to get:
6x(x-3) + 2(x^2-9) = x(5x-3)
----
6x^2-18x + 2x^2-18 = 5x^2 - 3x
------
8x^2 -18x - 18 = 5x^2 - 3x
------
3x^2 -15x - 18 = 0
------
x^2 - 5x - 6 = 0
----
(x-6)(x+1) = 0
x = 6 or x = -1
======================
Cheers,
Stan H.
======================